In 2013, 1000 patients were seen with a healthcare service with an average price of $50. this year, the company will pay half of the price. The demand for service is inelastic with elasticity coefficient of 0.8 between price $25 and $50. how many patients will take the service? explain by calculating percentage change in price to the patient
1
Expert's answer
2019-07-11T09:43:50-0400

elasticity coefficient =percentage change in quantity/percentage change in price

percentage change in price is: (25-50)/50 *100%= -50%

percentage change in quantity = percentage change in price*elasticity coefficient = 50*0,8 = 40%

percentage change in quantity is: (х-1000)/1000*100%= 40%

х - 1400 patients will take the service.